Landry Co. had the following information for the month of December. All direct materials were one hundred percent complete, aMultiple Choice $64.789. $168.820. $75.550. $80.402. $78.426.
Answer #1

Equivalent unit of material = (580*0%)+550+(445*100%) = 995 Unit

Cost of material = 80000

Cost per material = 80000/995 = $80.402

So answer is d) $80.402